Tohban Report 2014-08-13

From RHESSI Wiki

Jump to: navigation, search

Tohban Reports
Start Date: 6 Aug 2014
End Date: 13 Aug 2014
Tohban: Juan C. Martinez Oliveros
Tohban email:
Next Tohban: Albert Shih
List all reports


Solar Activity

Solar activity was very low this week. There are currently 7 ARs on the disk with a low probability to have a major flare in the following days.

How many GOES flares occurred?

 Flares above B, C, M, X class were     19    11     0     0

RHESSI flare list contains 0 over the time range 06-Aug-14 13-Aug-14

Memory Management

The SSR was filling up by the turn on procedure in the las two days. The SSR fill level reached a maximum of 79.2% and a minimum of 0% during the week.

Spacecraft Status

The cold plate 1 and 2 are at about 117.5 K and 115.9 K respectively.

Data Gaps

No data gaps

Detector issues

Other notes

Spacecraft Management

The cooldown procedure was completed on August 11 2014 and the operations team managed to get the cold plate back to basically the same temperature registered before the anneal. The cooler is running at ~67 W, which is as far as we can push it given the cold tip temperatures. The detectors were turned on the detectors also on August 11 2014 and started ramping the high voltages. On August 12 2014 we completed the high voltage ramps. Four of the detectors have segmented (1, 3, 8 and 9). The segmented detectors are currently at the nominal 4000 V. The voltage was increased on the five un-segmented detectors to 4500 V, but that was not enough to coerce any of them to segment.

The detectors were pretty noisy during the initial high voltage ramp, producing data faster than we could downlink during a pass. Some of this data was deleted, but not very much. The detectors settled down as the voltage was increased, especially those detectors which segmented. The un-segmented detectors required some tweaking of the thresholds, but are much better behaved now.

We re-enabled attenuator motions and both attenuators pulled out during the subsequent daylight portion of the orbit. We are currently using the following detectors for average livetime [1, 1, 3, 9].

Decimation Active/Vigorous
HLAT Decimation Rear decimation weight 6, no front decimation
Night time data (fronts) +-4 minutes
Night time data (rears) +-4 minutes
Require extra passes? Yes
Requirement for moving pointer? Yes
Attenuator operation Active
Detector problems? None
Personal tools